# Addressing the Rank Degeneration in Sequential Recommendation via Singular Spectrum Smoothing

### v1

@@ Abstract.
Sequential recommendation (SR) investigates the dynamic user preferences modeling and generates the next-item prediction. The next-item preference is typically generated by the affinity between the sequence and item representations. However, both sequence and item representations suffer from the rank degeneration issue due to the data sparsity problem. The representations with a degenerated rank become indistinguishable due to the lacking of informative components, and thus, weaken the recommendation performance. The rank degeneration issue significantly impairs the representations for SR. This motivates us to measure how severe is the rank degeneration issue and alleviate the sequence and item representation rank degeneration issues simultaneously for SR.
In this work, we theoretically connect the sequence representation degeneration issue with the item rank degeneration, particularly for short sequences and cold items. We also identify the connection between the fast singular value decay phenomenon and the rank collapse issue in transformer sequence output and item embeddings. We propose the area under the singular value curve metric to evaluate the severity of the singular value decay phenomenon and use it as an indicator of rank degeneration. We further introduce a novel singular spectrum smoothing regularization to alleviate the rank degeneration on both sequence and item sides, which is named after the Singular sPectrum sMoothing for sequential Recommendation (SPMRec). We also establish a correlation between the ranks of sequence and item embeddings and the rank of the user-item preference prediction matrix, which can affect recommendation diversity. We conduct experiments on four benchmark datasets to demonstrate the superiority of SPMRec over the state-of-the-art recommendation methods, especially in short sequences. Further analysis reveals that our proposed method mitigates the issue of fast singular value decay, thus preserving representation rank. The experiments also demonstrate a strong connection between our proposed singular spectrum smoothing and recommendation diversity. The code is in https://github.com/zfan20/SPMRec.
Sequential Recommendation, Degeneration, Rank, Spectrum, Nuclear Norm, Singular Values
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2023; Woodstock, NYbooktitle: Recsys ’23: ACM Conference on Recommender Systems, June 03–05, 2023, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
(a) SASRec User Sequence Output Embeddings
(b) SASRec Item Embeddings
Figure 1. The normalized singular values of user sequence output and item embeddings in the four benchmark categories of Amazon Review dataset. Singular values are normalized by dividing the largest singular value.
@@ 1. Introduction
Recommender systems play a significant and widespread role in various applications, as evident in numerous scenarios. Personalized recommendations have gained immense popularity, and sequential recommendation (SR) has recently garnered considerable attention due to its scalability and superior performance. SR utilizes sequential user behaviors and item-item transition correlations to predict the user’s next preferred item. Based on the underlying sequential model architectures, SR methods consist of Markov-chain based approaches (Rendle et al., 2010; He and McAuley, 2016), convolution-based approaches (Tang and Wang, 2018; Yan et al., 2019), RNN-based approaches (Hidasi et al., 2015; Zheng et al., 2019), and Transformer-based methods (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020). Among existing SR methods, Transformer-based methods have attracted increasing interest due to their capability of modeling high-order item-item transitions and scalable architecture design.
Despite the effectiveness of existing Transformer-based methods for SR, we argue that two rank degeneration issues exist in SR: (1) sequence embeddings suffer from the rank collapse issue and (2) items have the degeneration issue in the representation learning. We conducted an empirical analysis of singular values of sequence output and item embeddings in four benchmark Amazon Review datasets by SASRec (Kang and McAuley, 2018), which is shown in Fig. (1), where fast singular value decay in both sequence and item sides is observed. We conjecture that the fast singular value decay of representations of sequence outputs and items is a principal indicator of the rank degeneration issue observed in NLP application on word embedding (Wang et al., 2020). The fast singular value decay indicates that the largest singular value dominates and other singular values are much smaller, with the potential risks of losing the embeddings rank and embeddings being distributed as a narrow cone in the latent space. The rank losing implies the lack of independent and informative components to differentiate embeddings in the latent space. We argue that the distinct long-tail distributions of sequence length and item frequency in SR cause the rank degeneration issue for both sequence and item representations learning (Gao et al., 2019; Yu et al., 2022; Qiu et al., 2022). The failure to distinguish representations undermines the recommendation performance.
We argue that the item degeneration issue introduces sequence rank collapse, implying that sequence and item degeneration issues are intertwined. Several recent works (Dong et al., 2021; Zhai et al., 2023) discuss the exponential rank collapse of the Transformer due to the self-attention module if the residual connection is removed. However, it is still an open question to clearly pinpoint the cause of the Transformer rank collapse when we consider the Transformer as a whole. We theoretically demonstrate the likelihood of sequence and item rank collapse issue in Section 4.1, which will be more severe in short sequences and cold items modeling. An intuitive understanding is that the user-item preference, calculated as the affinity between the sequence and item representations, connects the deterioration of items with the sequence degeneration. We further disclose the connection between the ranks of sequence and item representations and the diversity of the user-item preference prediction matrix, with the theoretical (Section 4.4) and empirical analyses (Section 5.8). The intuition is that the rank of the user-item preference matrix correlates with the recommendation diversity and is determined by the ranks of sequence and item embeddings. We need a proper design to regularize the sequence and item representations learning to alleviate the rank degeneration.
@@
To this end, we propose a novel metric to measure the degree of rank degeneration issue and a novel singular spectrum smoothing regularization design to achieve better sequence and item representation learning. Specifically, as the fast singular value decay indicates the rank degeneration, we propose the Area Under the Singular value Curve (AUSC) to measure the degree of rank degeneration. Intuitively, a flat curve has a larger area under the curve than a fast decaying curve, which will also be illustrated in the following Fig. (2). We further propose the singular spectrum smoothing regularization loss as an auxiliary loss to constrain the sequence and item embedding learning, which is a surrogate of our proposed AUSC metric. We refer to the recommendation optimization with the proposed singular spectrum smoothing regularization as the Singular sPectrum sMoothing for sequential Recommendation (SPMRec). The singular spectrum smoothing regularization builds upon the nuclear norm and Frobenius norm (Horn and Johnson, 2012; Chen et al., 2022; Cui et al., 2020) to suppress the largest singular value and enlarge other singular values. Minimizing singular spectrum smoothing regularization alleviates the fast singular value decay effect and preserves the ranks of both sequence and item embeddings.
We summarize our five technical contributions as follows:
  • To the best of our knowledge, we are the first to investigate the sequence and item rank degeneration issues simultaneously in SR. We identify the connection between the fast singular value decay and the rank collapse in sequence and item embeddings and introduce a novel approach on moderating the singular value decaying phenomenon to alleviate rank collapse issues, hence avoiding embeddings being distributed in a narrow cone.
@@
  • We propose a novel metric (AUSC) to measure the degree of rank degeneration and develop a singular spectrum smoothing regularization named SPMRec with the nuclear norm and Frobenius norm to regularize both the sequence and item embeddings.
  • We theoretically and empirically demonstrate the existence of sequence and item rank collapse issues, which are more severe in short sequences and cold items learning.
  • Our experiments on four benchmark datasets show the effectiveness of the proposed singular spectrum smoothing regularization on alleviating the fast singular value decay and preserving ranks of sequence and item embeddings.
  • We connect the rank preserving of sequence and item embeddings with the rank of the user-item preference matrix, which strongly correlates with the recommendation diversity. Our experimental results show the connection and observe that a proper contribution of smoothing improves both recommendation performance and diversity while over-emphasis deteriorates the performance significantly.
2. Related Work
2.1. Sequential Recommendation
Sequential Recommendation (SR) is a technique that encodes the temporal interactions of each user as a sequence and uses the encoded behaviors to infer their preferences. The main idea behind SR is to model the transitions between items within sequences. The earliest SR methods used Markovian approaches, such as FPMC (Rendle et al., 2010) and Fossil (He and McAuley, 2016). FPMC assumes that the item prediction depends only on the most recent previous interacted item, which is the first-order Markovian assumption. Fossil extends FPMC with the addition of item similarities information. Another line of work utilizes the Recurrent Neural Network (RNN) as the base model for modeling the temporal order of sequential interactions. The recurrent modeling characteristics of RNN make it suitable for modeling temporally ordered behaviors. Several works adopt RNNs to address the SR problem, including GRU4Rec (Hidasi et al., 2015), HGN (Ma et al., 2019; Peng et al., 2021), HRNN (Quadrana et al., 2017), and RRN (Wu et al., 2017). Different variants of RNNs are proposed, such as the original GRU in GRU4Rec and hierarchical RNN in HGN and HRNN. The third line of methods to solve the SR problem is using the Convolution Neural Network (CNN). With the interpretation of a sequence of item embeddings as an image, convolution operators are introduced to extract long-term and short-term user dynamic interests. The representative works include Caser (Tang and Wang, 2018), CosRec (Yan et al., 2019), and NextItNet (Yuan et al., 2019). Caser proposes convolutions to latent dimensions and item perspectives for comprehensive sequence encoding. Instead of treating the sequence as an image, CosRec extends to represent the sequence as a 3D tensor and adopts the 2D convolution operators. The NextItNet proposes the dilated convolution in the sequence modeling. The recent success of the self-attention module (Vaswani et al., 2017) inspires the adaptation of Transformer to the sequential recommendation task. The self-attention module captures the sequential high-order item transitions by modeling attention from all items in the sequence. SASRec (Kang and McAuley, 2018) is the earliest work of applying Transformer to SR. BERT4Rec (Sun et al., 2019) builds upon BERT (Devlin et al., 2019) with bidirectional Transformers for SR. FMLP-Rec (Zhou et al., 2022) removes the self-attention module in the Transformer and proposes an all-MLP architecture for SR. Several variants of Transformers (Fan et al., 2021; Li et al., 2020; Wu et al., 2020; Liu et al., 2021) are proposed to enhance the SR further.
2.2. Representation Degeneration
The representation degeneration problem describes the issue of learned embeddings being narrowly distributed in a small cone (Gao et al., 2019). The degeneration issue is firstly proposed in the word representations learning problem, where word embeddings are randomly initialized and then optimized via gradient descent. As shown in (Gao et al., 2019), words from text data follow the long-tail distribution, where most words have a low frequency of occurrence in the data. The long-tail distribution of words degenerates the representation learning process. Theoretically, it causes the narrow distribution of words in the representation latent space (Gao et al., 2019). Moreover, an empirical analysis (Yu et al., 2022) shows that the learning of tail words degenerates the representation optimization of frequent words. An important analysis in (Wang et al., 2020) shows that degeneration is closely related to the fast decaying singular values distribution. In other words, the largest singular value is significantly larger than all other singular values.
Distinct from the natural language processing (NLP) area, SR still has several unique characteristics. For example, sequences are short (data sparsity in recommendation systems), while sentences/paragraphs in NLP are typically long. This distinct characteristic introduces additional challenges for SR, where the degeneration issue might also happen in the sequence output embeddings. DuoRec (Qiu et al., 2022) observes that the representation degeneration issue also exists in the self-supervised SR, with data augmentations. DuoRec and DirectAU (Wang et al., 2022) optimizes the uniformity property in the in-batch representations to enforce the representations to be uniformly distributed, where the uniformity is based on the Euclidean distance. This work connects the item embeddings degeneration with the degeneration of sequence representations. The intertwining of both sequence and item rank degeneration is rarely discussed before in SR. Moreover, some solutions from NLP to the degeneration problem only apply to the item side, including the CosReg (Gao et al., 2019) based on cosine similarities, Euclidean regularization based on learned word similarities (Zhang et al., 2020), spectrum control (Wang et al., 2020; Yan et al., 2022), and gradient gating (Yu et al., 2022). The general idea of these works is regularizing the representation matrix by strengthening the differences between representations, mostly either based on cosine similarities (e.g., CosReg (Gao et al., 2019)) or Euclidean methods (e.g., (Zhang et al., 2020) and the uniformity property adopted by DuoRec and DirectAU). Unlike existing solutions, this work investigates the rank degeneration issue from the singular value perspective of representations and proposes a novel regularization, which preserves the rank of representations and keeps representations distinguishable.
@@ 3. Preliminaries
This section presents the formal definition of the SR problem and the adoption of the Transformer in SR.
@@ 3.1. Problem Definition
In the SR problem, there is a set of users $\mathcal{U}$, a set of items $\mathcal{V}$, and interactions feedback between users and items. Each interaction is associated with its timestamp. For each user $u\in\mathcal{U}$, we can sort the interacted items chronologically and format the user as the sequence of interacted items, denoted as $\mathcal{S}^{u}={[v^{u}_{1},v^{u}_{2},\dots,v^{u}_{|\mathcal{S}^{u}|}]}$. Specifically, $v^{u}_{i}\in\mathcal{V}$ represents the $i$-th interacted item in the sequence $\mathcal{S}^{u}$ of the user $u$. SR addresses the next item prediction problem given the user history interaction sequence, which is formalized as predicting the probability of the next item $p\left(v_{|\mathcal{S}^{u}|+1}^{(u)}=v\left|\mathcal{S}^{u}\right.\right)$. In the recommendation generation stage, the next item prediction scores over all items are ranked to generate the top-N recommendation list for the user $u$.
3.2. Transformer for SR
The fundamental challenge in SR is the sequential modeling of users’ dynamic behaviors. Among existing sequential modeling methods for SR, Transformer-based methods (Kang and McAuley, 2018; Sun et al., 2019) demonstrate superior effectiveness in capturing dynamic interests and modeling different orders of item-item transitions. Following the widely adopted Transformer architecture (Vaswani et al., 2017), Transformer-based SR methods consist of several essential components, including residual connection, layer normalization, self-attention module, and feed-forward networks.
Specifically, for a given user interaction sequence $\mathcal{S}^{u}$, we pre-process it by either truncating earliest interactions or padding with zeros such that the sequence length becomes the maximum sequence length $n$ to get a fixed length sequence $s=(s_{1},s_{2},\dots,s_{n})$. An item embedding matrix $\mathbf{M}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is randomly initialized and will be optimized, where $d$ is the latent dimension size. A trainable positional embedding $\mathbf{P}\in\mathbb{R}^{n\times d}$ is added to the sequentially ordered item embedding matrix as:
(1) $\mathbf{E}_{\mathcal{S}^{u}}=[\mathbf{m}_{s_{1}}+\mathbf{p}_{s_{1}},\mathbf{m}_{s_{2}}+\mathbf{p}_{s_{2}},\dots,\mathbf{m}_{s_{n}}+\mathbf{p}_{s_{n}}].$
Following the original Transformer (Vaswani et al., 2017), existing Transformer-based sequential methods include Feed-forward Neural Networks, Residual Connection, and Layer Normalization modules. Formally, we formulate the sequence encoder as:
@@ 3.2. Transformer for SR
where $\mathbf{H}_{u}\in\mathbb{R}^{n\times d}$ denotes sequential output embeddings of $\mathcal{S}^{u}$, and for each timestep $t$, and $\textbf{H}_{(u,t)}$ encodes the predicted next-item ($v_{t+1}$) representation. The next item preference score prediction is given as:
@@
where $\mathbf{E}_{v}$ denotes the embedding of the item $v$. The typical optimization process adopts the sampled cross-entropy loss or bayesian personalized ranking loss (BPR) (Rendle et al., 2012), which is maximizing the probability of users ranking positive items $v^{+}_{t+1}$ with greater preference scores over sampled negative items $v^{-}_{t+1}$. Specifically, the maximization of probabilistic likelihood over all users and all time steps is as follows:
(4) $\text{max}\sum_{\mathcal{S}^{u}\in\mathcal{S}}\sum_{t=1}^{|\mathcal{S}^{u}|}p\left(v^{+}_{t+1}>v^{-}_{t+1}\left|\mathcal{S}^{u}\right.\right),$
where $v^{+}_{t+1}>v^{-}_{t+1}$ means that the positive item $v^{+}_{t+1}$ is ranked ahead of the negative item $v^{-}_{t+1}$, which can be further reformulated as a sampled cross-entropy loss minimization as:
(5) $\mathcal{L}_{\text{rec}}=\sum_{\mathcal{S}^{u}\in\mathcal{S}}\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\frac{\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v^{+}_{t+1}}\right)}{\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v^{+}_{t+1}}\right)+\sum_{j=1}^{N^{-}_{u}}\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v^{-}_{t+1}}^{j}\right)},$

### v2

# Sequential Recommendation with Controllable Diversification: Representation Degeneration and Diversity

@@ Abstract.
Sequential recommendation (SR) models the dynamic user preferences and generates the next-item prediction as the affinity between the sequence and items, in a joint latent space with low dimensions (i.e., the sequence and item embedding space). Both sequence and item representations suffer from the representation degeneration issue due to the user/item long-tail distributions, where tail users/items are indistinguishably distributed as a narrow cone in the latent space. We argue that the representation degeneration issue is the root cause of insufficient recommendation diversity in existing SR methods, impairing the user potential exploration and further worsening the echo chamber issue.
In this work, we first disclose the connection between the representation degeneration and recommendation diversity, in which severer representation degeneration indicates lower recommendation diversity. We then propose a novel Singular sPectrum sMoothing regularization for Recommendation (SPMRec), which acts as a controllable surrogate to alleviate the degeneration and achieve the balance between recommendation diversity and performance. The proposed smoothing regularization alleviates the degeneration by maximizing the area under the singular value curve, which is also the diversity surrogate. We conduct experiments on four benchmark datasets to demonstrate the superiority of SPMRec, and show that the proposed singular spectrum smoothing can control the balance of recommendation performance and diversity simultaneously.
Recommendation, Nuclear Norm, Determinant, Diversity
(a) Beauty
(b) Toys
Figure 1. A demonstrated strong correlation between the proposed spectrum smoothing regularization (weighted by $\beta$) and intra-list diversity on public Amazon review data. A challenging balance between recommendation performance and diversity can be achieved with the proposed method.
@@ 1. Introduction
Recommender systems play a significant and widespread role in providing personalized recommendations and user modeling for various web applications, as evident in fashion, music, movies, and marketing scenarios. Among existing frameworks, sequential recommendation (SR) (Rendle et al., 2010; He and McAuley, 2016; Tang and Wang, 2018; Hidasi et al., 2015; Kang and McAuley, 2018) has recently garnered considerable attention due to its scalability and superior performance. SR utilizes sequential user behaviors and item-item transition correlations to predict the user’s next preferred item. Among existing SR methods, Transformer-based methods (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020) have attracted increasing interests due to their capability of modeling high-order item-item transitions and scalable architecture design. In addition to recommendation quality (quantified as next-item prediction accuracy), a diverse set of items generation is also crucial for improving user satisfaction (Zhang and Hurley, 2008; Chen et al., 2018) and helping users escape the echo chamber (Ge et al., 2020) for better long-term engagements (Chen et al., 2021).
The long-tail distribution of users/items has been widely observed in real-world applications and benchmarking datasets (Zhang et al., 2023; Domingues et al., 2012; Wang et al., 2022a), in which a small portion of users/items contribute to most interactions. According to empirical observations (Goel et al., 2010), tail items availability in recommendation can boost the sales of head items, which triggers the need of diverse recommendations. Demonstrated by (Wang et al., 2018; Gao et al., 2023), the long-tail distribution potentially leads to more intense Matthew effect, in which popular items and categories further dominate the recommender system with the implication of insufficient recommendation diversity. However, few research investigates the relationship between the long-tail distribution of users/items and the recommendation diversity.
Existing methods either focus on post re-ranking approaches (with the risk of sacrificing recommendation quality) (Zhang and Hurley, 2008; Chen et al., 2018; Wilhelm et al., 2018) or build upon the indirect connection between recommendation quality and diversity (Yang et al., 2023; Liu et al., 2022). A balance between recommendation quality and diversity is needed because over-emphasizing either the quality or the diversity leads to sub-optimal user satisfaction. To achieve this balance, a recommendation model with controllable diversification module is needed. It is also challenging to develop the controllable diversity module because the underlying connection between the quality and the diversity is not identified yet, and a differentiable solution is preferred so as to optimize both the quality and diversity.
(a) User Sequence Embeddings
(b) Item Embeddings
Figure 2. Fast singular values decay of user sequence output and item embeddings of SASRec are observed, implying the representation degeneration exists in both users and items. Singular values are normalized by dividing the largest singular value.
To this end, we demonstrate the connection between the long-tail distribution on users/items and the recommendation diversity, via the observed representation degeneration issue in both users and items, as shown in Fig. (2). We theoretically show that the fast singular value decay from the users/items representation degeneration is the root cause of insufficient diversity. More details will be explained in Section  3. Motivated by this connection, we develop a novel sequential recommendation with controllable diversity module by smoothing the singular value curves for users/items representations, with the goals of alleviating the degeneration and improving the recommendation diversity simultaneously.
@@
Specifically, we propose a novel metric to measure the degree of degeneration issue and a novel Singular sPectrum sMoothing regularization design to achieve better sequence and item representation learning for Recommendations, SPMRec. We propose the Area Under the Singular value Curve (AUSC) to measure the degree of degeneration, with the intution that a flat curve has a larger area under the curve than a fast decaying curve, which will also be illustrated in the following Fig. (3). We further propose the singular spectrum smoothing regularization loss as an auxiliary loss to constrain the sequence and item embedding learning, which is a surrogate of our proposed AUSC metric. The singular spectrum smoothing regularization builds upon the nuclear norm and Frobenius norm (Horn and Johnson, 2012; Chen et al., 2022; Cui et al., 2020) to suppress the largest singular value and enlarge other singular values. The singular spectrum smoothing regularization unblocks the feasibility of sequential recommendation with controllable diversity. We summarize our three technical contributions as follows:
  • To the best of our knowledge, we are the first to identify the connection between the representation degeneration and insufficient diversity in SR from the theoretical perspective.
@@
  • We propose a novel metric (AUSC) to measure the degree of degeneration and develop a singular spectrum smoothing regularization named SPMRec to achieve the important but challenging balance of recommendation quality and diversity.
  • Our experiments on four benchmark datasets show the effectiveness of the proposed SPMRec, and empirically demonstrate that a proper contribution of smoothing improves both recommendation performance and diversity.
@@ 2.1. Problem Definition
In the SR problem, there are a set of users $\mathcal{U}$, a set of items $\mathcal{V}$, and interactions feedback between users and items. Each interaction is associated with its timestamp. For each user $u\in\mathcal{U}$, we sort the interacted items chronologically and format the user as the sequence of interacted items, denoted as $\mathcal{S}^{u}={[v^{u}_{1},v^{u}_{2},\dots,v^{u}_{|\mathcal{S}^{u}|}]}$. Specifically, $v^{u}_{i}\in\mathcal{V}$ represents the $i$-th interacted item in the sequence $\mathcal{S}^{u}$ of the user $u$. SR tackles the problem of predicting the next item based on the sequence of user’s historical interactions, which is formalized as predicting the probability of the next item $p\left(v_{|\mathcal{S}^{u}|+1}^{(u)}=v\left|\mathcal{S}^{u}\right.\right)$. In the recommendation generation stage, the next item prediction scores over all items are ranked to generate the top-N recommendation list for the user $u$.
2.2. Sequential User Modeling
Given a user interaction sequence $\mathcal{S}^{u}$, we pre-process it by either truncating earliest interactions or padding with zeros such that the sequence length becomes the maximum sequence length $n$ to get a fixed length sequence $s=(s_{1},s_{2},\dots,s_{n})$. An item embedding matrix $\mathbf{M}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is randomly initialized and will be optimized, where $d$ is the latent dimension size. A trainable positional embedding $\mathbf{P}\in\mathbb{R}^{n\times d}$ is added to the sequentially ordered item embedding matrix as $\mathbf{E}_{\mathcal{S}^{u}}$. Formally, we formulate the sequence encoder as follows:
@@ 2.2. Sequential User Modeling
where $\mathbf{H}_{u}\in\mathbb{R}^{n\times d}$ denotes sequential output embeddings of $\mathcal{S}^{u}$, and for each timestep $t$, and $\textbf{H}_{(u,t)}$ encodes the predicted next-item ($v_{t+1}$) representation.
2.3. Next-Item Prediction Optimization
The next item preference score prediction is given as:
@@ 2.3. Next-Item Prediction Optimization
where $\mathbf{E}_{v}$ denotes the embedding of the item $v$. The typical optimization process adopts the classical sampled cross-entropy (sampled CE) loss or bayesian personalized ranking loss (BPR) (Rendle et al., 2012), which is maximizing the probability of users ranking positive items $v^{+}_{t+1}$ with greater preference scores over sampled negative items $v^{-}_{t+1}$. The optimization can be formulated as a sampled cross-entropy loss minimization as:
(3) $\mathcal{L}_{\text{rec}}=\sum_{\mathcal{S}^{u}\in\mathcal{S}}\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\frac{\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{M}_{v^{+}_{t+1}}\right)}{\sum_{j=1}^{N^{-}_{u}}\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{M}_{v^{-}_{t+1}}^{j}\right)},$